One Guava has a mass of 2.0 grams and one Jelly has a mass of 1.5 g. How many Guava Jelly can be made with 12.5 grams of Guava and 15.0 grams of Jelly? (Show your solutions)
1
Expert's answer
2021-10-20T02:52:02-0400

1 Gu + 4 Je → 1 GuJe

We have 12.5g/2.0g = 6.25 Guavas and 15.0g/1.5g = 10 Jellys. From it we can make 10/4 = 2.5 Guava Jellys and have 6.25 - 2.5 = 3.75 Guavas left.


Answer: 2 Guava Jellys (rounded to the nearest integer).